OpenStudy (anonymous):

Find the domain of the given function. f(x) = sqrt. x+6/{(x+7)(x-9)}

OpenStudy (anonymous):

Is your question to find the domain of \[\sqrt{\frac{ x+6 }{ (x+7)(x-9) }}\]

OpenStudy (anonymous):

no the sqrt is only on the top @calmat01

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So, you know that you cannot take square roots of negative numbers, correct?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

And you also know that you cannot divide by zero, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Alright, so now, all we have to do is make sure the x's we use satisfy the two situations. Namely, x+6>=0 and also x+7 cannot be zero nor can x-9 by zero.

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

so x can not be = to 7,6, and 9?

OpenStudy (anonymous):

Not quite, but close. Since \[x+6\ge0\] because we cannot take the sq. rt of a negative number, the first thing we know is that x must be greater than or equal to -6. Now since x+7 =0 when x=-7, I can't use that value because we have already established that x must be greater than or equal to -6. Finally, x-9=0 when x=9, so you can use any number less than 9 but greater than or equal to -6, and you can use any number greater than 9. Put all of that together, and your domain is \[[-6,9)\cup\left( 9, \right)\]